OpenStudy (anonymous):

solve the equation log base 3 (2x-1)=2

OpenStudy (anonymous):

\[\log_{3}(2x-1)=2 \\ can \ be \ rewritten \ as \\ 3^2 = 2x-1\]

OpenStudy (anonymous):

how do you solve it?

OpenStudy (fibonaccichick666):

how dlouw ouy isolate the x

OpenStudy (fibonaccichick666):

how would you isolate the x*

OpenStudy (anonymous):

make it a neg number?

OpenStudy (fibonaccichick666):

please show your idea here

OpenStudy (anonymous):

-2

OpenStudy (fibonaccichick666):

ok how about if I gave you 8x=4 can you please solve this?

OpenStudy (anonymous):

nope

OpenStudy (fibonaccichick666):

then if you won't attempt it, I will not help you

OpenStudy (anonymous):

i can't attempt if I dont know what your doing

OpenStudy (fibonaccichick666):

solve: 8x=4 please

OpenStudy (anonymous):

=1/2

OpenStudy (fibonaccichick666):

ok now how did you solve it?

OpenStudy (anonymous):

I divided

OpenStudy (fibonaccichick666):

please show me your steps no matter how trivial

OpenStudy (anonymous):